Mr. Wilson invested money in two accounts. His total investment was $20,000. If one account pays 2% in interest and the other pays 5% in interest, how much does he have in each account if he earned a total of $550 in interest in 1 year?
In the 2% account: $
In the 5% account: $

Answer by mananth(16949)

part I 2.00% per annum ---x
part II 5.00% per annum ---y

x + 1 y= 20000 ------------------------1
2.00% x + 5.00% y= 550
Multiply by 100
2 x + 5 y= = 55000.00 --------2

Multiply (1) by -2
we get

-2 x -2 y= -40000.00

Add this to (2)

0 x 3 y= = 15000

divide by 3

y = 5000 investment at 5.00%
Balance 15000 investment at 2.00%
CHECK
5000 --------- 5.00% ------- 250.00
15000 ------- 2.00% ------- 300.00
Total -------------- 550.00